if you collect 1.75 L of hydrogen gas during a lab experiment, when the room temperature is 23ºC  and the barometric pressure is 105 kPa, how many moles of hydrogen will you have?


1
Expert's answer
2021-02-01T03:52:14-0500

using gas law,

"P*V =n *R*T"

Using given data,

n "=\frac{P*V}{R*T}"

therefore, n = 0.074 moles